# Stochastic Cahn–Hilliard Equations from One-Dimensional Ising–Kac–Kawasaki Dynamics

### 31 Mar 2026 v1

@@ 2.2 Well-posedness of the 11d stochastic Cahn–Hilliard equation
The shifted equation for $Y$ becomes a random but classical semilinear fourth-order parabolic equation, and standard $H^{-1}$-energy estimates in $H^{-1}$, $H^{1}$, and $L^{4}$ become available. Testing (6) against $(-\partial_{x}^{2})^{-1}Y$ yields the formal coercive identity
$\frac{1}{2}\frac{d}{dt}\|Y(t)\|_{H^{-1}}^{2}+\nu\|Y(t)\|_{H^{1}}^{2}$ $=-\lambda A\,\|Y(t)\|_{L^{2}}^{2}-\lambda\chi\int_{\mathbb{T}}Y(t,x)\,(Y(t,x)+Z(t,x))^{3}\,dx.$ (8)
@@
We summary the well-posedness of the one-dimensional stochastic Cahn–Hilliard equation as follows.
@@ 2.3 Dirichlet forms, entropy inequality, and the Kipnis–Varadhan bound
In this subsection we collect the canonical Gibbs measures naturally associated with the conservative Kawasaki dynamics, both on the full torus and on finite blocks, together with the corresponding Dirichlet forms and Fisher information. Standard references are Guo–Papanicolaou–Varadhan [12], Yau [22], Kipnis–Varadhan [16], and Kipnis–Landim [15].
On each invariant sector $\Sigma_{N,M}$, the canonical Gibbs measure $\mu_{N,\gamma,\beta}$ is the natural reversible reference measure.
Definition 2.1 (Global Dirichlet form on the canonical sector).
@@
$D_{N,\gamma}(f):=-\bigl\langle f,\mathscr{L}_{\gamma}^{N}f\bigr\rangle_{\mu_{N,\gamma,\beta}^{M}}=N^{2}\sum_{i\in\Lambda_{N}}D_{i}^{N,\gamma},$
@@
The corresponding global Fisher information is $I_{N,\gamma}(f):=D_{i}^{N,\gamma}(\sqrt{f}),$ for every density $f\geq 0$ with $\int_{\Sigma_{N,M}}f\,d\mu_{N,\gamma,\beta}^{M}=1$.
Fix $\ell\geq 1$, and let $\Lambda_{\ell}:=\{-\ell,\dots,\ell\}$. The block Hamiltonian is
$H_{\gamma,\ell}^{\mathrm{blk}}(\eta):=-\frac{1}{2}\sum_{u,v\in\Lambda_{\ell}}\kappa_{\gamma}(u-v)\eta_{u}\eta_{v},\qquad\eta\in\sum_{\ell}:=\{-1,1\}^{\Lambda_{\ell}}.$
@@
where the normalizing constant $Z_{\gamma,\ell,m}^{\mathrm{can}}:=\sum_{\eta\in\Sigma_{\ell,m}^{\mathrm{can}}}e^{-\beta H_{\gamma,\ell}^{\mathrm{blk}}(\eta)}.$ We remark that the grand canonical block Gibbs measure can be obtain by linear combination of canonical block Gibbs measures.
@@
Since the exchange $\eta\mapsto\eta^{j,j+1}$ only modifies two spins, the corresponding energy increment
$\Delta_{j}^{(\ell)}H_{\gamma,\ell}^{\mathrm{blk}}(\eta)$
is uniformly bounded for fixed $\ell$, because $\sum_{z}\kappa_{\gamma}(z)=1$ and $|\eta_{u}|\leq 1$. The logistic form of the rate therefore yields uniform upper and lower bounds on $c_{j,j+1}^{(\ell),\gamma}(\eta)$. The same argument applies to the full-torus rates $c_{i}^{\gamma}(\sigma)$, since $|\Delta_{i}H_{\gamma}(\sigma)|$ is uniformly bounded by the $L^{1}$-normalization of the kernel. ∎
@@
The first inequality is the standard entropy inequality; see [15, Lemma 6.1.4] or [12, Section 2]. Applying it to $F=|Y|$ and using
@@
Note that the Markov generator $\mathscr{L}_{\gamma}^{N}$ is a self-adjoint non-positive operator on $L^{2}(\mu_{N,\gamma,\beta})$. We recall the standard Kipnis–Varadhan estimate for additive functionals of reversible Markov processes in the following lemma; see [16] or [15, Chapter 2, Section 1].
@@
Now we show the entropy–Kipnis–Varadhan estimate for non-equilibrium initial laws
@@
Let $f_{N,0}\mu_{N,\gamma,\beta}$ be the initial law, and assume $H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})\leq C_{H}N.$ Let $W:\Sigma_{N}\to\mathbb{R}$ be bounded with $\int W\,d\mu_{N,\gamma,\beta}=0.$ Under the entropy bound $H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})\leq C_{H}N$, we have
$\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}W(\sigma(t))\,dt\right|\leq C\,\sqrt{NT}\,\|W\|_{\mathcal{H}^{-1}_{N}}.$ (11)
@@
Lemma 6 with $Y:=\int_{0}^{T}W(\sigma(t))\,dt,\mu=f_{N,0}\mu_{N,\gamma,\beta}$, and $\nu=\mu_{N,\gamma,\beta}$ yields that for every $a>0$,
$\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}W(\sigma(t))\,dt\right|\leq\frac{H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})+\log 2}{a}+\frac{1}{a}\max_{\pm}\log\mathbb{E}_{\mu_{N,\gamma,\beta}}\left[\exp\!\left(\pm a\int_{0}^{T}W(\sigma(t))\,dt\right)\right].$
This gives
$\mathbb{E}_{f_{N,0}}|Y|\leq\frac{H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})+\log 2}{a}+\frac{1}{a}\max_{\pm}\log\mathbb{E}_{\mu_{N,\gamma,\beta}}\bigl[e^{\pm aY}\bigr].$
@@
$\log\mathbb{E}_{\mu_{N,\gamma,\beta}}\left[\exp\!\left(\pm a\int_{0}^{T}W(\sigma(t))\,dt\right)\right]\leq C_{\mathrm{KV}}\,a^{2}T\,\|W\|_{\mathcal{H}^{-1}_{N}}^{2}.$
Substitution gives
$\mathbb{E}_{f_{N,0}}|Y|\leq$ $\frac{H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})+\log 2}{a}+C_{\mathrm{KV}}aT\|W\|_{\mathcal{H}^{-1}_{N}}^{2}$
$\leq$ $C\sqrt{\bigl(H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})+1\bigr)\,T}\,\|W\|_{\mathcal{H}^{-1}_{N}}.$
@@ 3 Expansion for the discrete drift
In this section we give a expansion for the discrete drift for the rescaled field $X_{\gamma}$.
Since the Ising–Kac Hamiltonian
$H_{\gamma}(\sigma)=-\frac{1}{2}\sum_{x,y\in\Lambda_{N}}\kappa_{\gamma}(x-y)\sigma_{x}\sigma_{y}=-\frac{1}{2}\sum_{y\in\Lambda_{N}}h(\sigma,y)\sigma_{y},$
@@
$X_{\gamma}(t,x)=X_{\gamma}(0,x)-\int_{0}^{t}\nabla_{\varepsilon_{\gamma}}\mathcal{J}_{\gamma}(s,x)\,ds+M_{\gamma}(t,x),$
@@
$\Bigl(\alpha_{\gamma}^{-1}\mathscr{L}_{\gamma}^{N}\langle X_{\gamma},\phi\rangle_{\gamma}\Bigr)(\sigma)=-\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}\mathcal{J}_{\gamma}(\sigma;\varepsilon_{\gamma}i)\nabla_{\varepsilon_{\gamma}}\phi(\varepsilon_{\gamma}i)=-\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}j_{i}(\sigma)\,\nabla_{\varepsilon_{\gamma}}\bigl(\mathcal{K}_{\gamma}^{\varepsilon}\phi\bigr)(\varepsilon_{\gamma}i)$
@@
$(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i):=\varepsilon_{\gamma}\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(j-i)\phi(\varepsilon_{\gamma}j)=\varepsilon_{\gamma}\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(i-j)\phi(\varepsilon_{\gamma}j).$
Since $X_{\gamma}(t,\varepsilon_{\gamma}i)=\delta_{\gamma}^{-1}h_{\gamma}\!\left(\sigma_{t/\alpha_{\gamma}};i\right),\varepsilon_{\gamma}i\in\mathbb{T}$, we have
$h_{\gamma}\!\left(\sigma\!\left(\frac{t}{\alpha_{\gamma}}\right);i\right)-h_{\gamma}\!\left(\sigma\!\left(\frac{t}{\alpha_{\gamma}}\right);i+1\right)=-\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i),$
@@
with $\widetilde{R}_{i}=\frac{1}{6}F^{(3)}(a_{i}+\theta_{i}b_{i})b_{i}^{3},\ \theta_{i}\in(0,1).$ Therefore
@@
(i) Linear part.
Since $d_{i}^{2}\in\{0,4\}$, we have $a_{i}^{\gamma}\in\{0,4\kappa_{\gamma}(1)\}$, hence $F(a_{i}^{\gamma})$ takes only two values. Using the Taylor expansion $F(u)=\frac{1}{2}-\frac{u}{4}+O(u^{2})$ at $u=0$, one obtains
@@
(ii) Nonlinear part.
Since $F^{\prime}(u)=-\frac{1}{4}+O(u)$, and $d_{i}b_{i}^{\gamma}=-d_{i}^{2}\,\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}$, we obtain the leading nonlinear contribution
@@
This is the term whose second-order Boltzmann–Gibbs closure produces the cubic drift in the limit.
(iii) Remainder.
By construction, $j_{i}^{\mathrm{rem}}=d_{i}R_{i}^{\gamma},R_{i}^{\gamma}=\frac{1}{2}F^{\prime\prime}(\theta_{i}^{\gamma})\,(b_{i}^{\gamma})^{2},$ and $F^{\prime\prime}$ is bounded on bounded sets. Hence, using $|d_{i}|\leq 2$ and $|b_{i}^{\gamma}|\lesssim\delta_{\gamma}\varepsilon_{\gamma}|\nabla_{\varepsilon_{\gamma}}X_{\gamma}|$, we have
$|j_{i}^{\mathrm{rem}}|\lesssim\delta_{\gamma}^{2}\varepsilon_{\gamma}^{2}\bigl|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(\varepsilon_{\gamma}i)\bigr|^{2}.$
In particular, under the same a priori discrete $H^{1}$-type bounds used in the tightness/drift analysis, the contribution of $j_{i}^{\mathrm{rem}}$ to the weak drift is negligible.
@@
$\mathcal{J}_{\gamma}^{\mathrm{lin}}(t,x)$ $:=\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}\kappa_{\gamma}\!\left(\frac{x}{\varepsilon_{\gamma}}-i\right)j_{i}^{\mathrm{lin}}\left(\sigma\!\left(\frac{t}{\alpha_{\gamma}}\right)\right)$ (13)
@@
$\mathcal{J}_{\gamma}^{\mathrm{nl}}(t,x)$ $:=\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}\kappa_{\gamma}\!\left(\frac{x}{\varepsilon_{\gamma}}-i\right)\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\,j_{i}^{\mathrm{nl}}\left(\sigma\!\left(\frac{t}{\alpha_{\gamma}}\right)\right),$ (14)
@@
$\mathcal{J}_{\gamma}^{\mathrm{rem}}(t,x)$ $:=\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}\kappa_{\gamma}\!\left(\frac{x}{\varepsilon_{\gamma}}-i\right)j_{i}^{\mathrm{rem}}.$ (15)
Here

### 05 Apr 2026 v2

@@ 2.2 Well-posedness of the 11d stochastic Cahn–Hilliard equation
The shifted equation for $Y$ becomes a random but classical semilinear fourth-order parabolic equation, and standard $H^{-1}$-energy estimates in $H^{-1}$, $H^{1}$, and $L^{4}$ become available. Testing (7) against $(-\Delta)^{-1}Y$ yields the formal coercive identity
$\frac{1}{2}\frac{d}{dt}\|Y(t)\|_{H^{-1}}^{2}+\nu\|Y(t)\|_{H^{1}}^{2}$ $=-A\,\|Y(t)\|_{L^{2}}^{2}-\chi\int_{\mathbb{T}}Y(t,x)\,(Y(t,x)+Z(t,x))^{3}\,dx.$ (9)
@@
We summarize the well-posedness of the one-dimensional stochastic Cahn–Hilliard equation as follows.
@@ 2.3 Dirichlet forms, entropy inequality, and the Kipnis–Varadhan bound
On each invariant sector $\Sigma_{N,M}$, the canonical Gibbs measure $\mu_{N,\gamma,\beta}$ is the natural reversible reference measure. In this subsection we collect the canonical Gibbs measures naturally associated with the conservative Kawasaki dynamics, both on the full torus and on finite blocks, together with the corresponding Dirichlet forms and Fisher information. Standard references are Guo–Papanicolaou–Varadhan [14], Yau [24], Kipnis–Varadhan [18], and Kipnis–Landim [17].
Definition 2.1 (Dirichlet form on the canonical sector).
@@
$D_{N,\gamma}(f):=-\bigl\langle f,\mathscr{L}_{\gamma}^{N}f\bigr\rangle_{\mu_{N,\gamma,\beta}}=\sum_{i\in\Lambda_{N}}D_{i}^{N,\gamma},$
@@
Using the Dirichlet form, we can define the Fisher information of the Ising–Kac–Kawasaki dynamics.
Definition 2.2.
For a density $f$ with respect to $\mu_{N,\gamma,\beta}$, we define the bond Fisher information by
$I_{i,i+1}^{N,\gamma}(f):=D_{j}^{N,\gamma}(\sqrt{f})=\frac{1}{4}\int_{\sum_{N,M}}c_{j}^{\gamma}(\sigma)\bigl(\sqrt{f(\sigma^{i,i+1})}-\sqrt{f(\sigma)}\bigr)^{2}\,\mu_{N,\gamma,\beta}(d\sigma).$
The corresponding global Fisher information is
$I_{N,\gamma}(f):=\sum_{i\in\Lambda_{N}}I_{i,i+1}^{N,\gamma}(f)=\sum_{i\in\Lambda_{N}}D_{i}^{N,\gamma}(\sqrt{f}),$
for every density $f\geq 0$ with $\int_{\Sigma_{N,M}}f\,d\mu_{N,\gamma,\beta}^{M}=1$.
Fix $\ell\geq 1$, and let $\Lambda_{\ell}:=\{-\ell,\dots,\ell\}$. Write a global configuration as $\sigma=(\eta,\xi)$, where $\eta\in\sum_{\ell}:=\{-1,1\}^{\Lambda_{\ell}}$ denotes the block variables, and $\xi$ the exterior variables. The block Hamiltonian is
$H_{\gamma,\ell}^{\mathrm{blk}}(\eta):=-\frac{1}{2}\sum_{u,v\in\Lambda_{\ell}}\kappa_{\gamma}(u-v)\eta_{u}\eta_{v}.$
@@
where the normalizing constant $Z_{\gamma,\ell,m}^{\mathrm{can}}:=\sum_{\eta\in\Sigma_{\ell,m}^{\mathrm{can}}}e^{-\beta H_{\gamma,\ell}^{\mathrm{blk}}(\eta)}.$ We remark that the grand canonical block Gibbs measure can be obtained by linear combination of canonical block Gibbs measures.
@@
Since the exchange $\eta\mapsto\eta^{j,j+1}$ only modifies two spins, the corresponding energy increment $\Delta_{j}^{(\ell)}H_{\gamma,\ell}^{\mathrm{blk}}(\eta)$ is uniformly bounded for fixed $\ell$, because $\sum_{z}\kappa_{\gamma}(z)=1$ and $|\eta_{u}|\leq 1$. The logistic form of the rate therefore yields uniform upper and lower bounds on $c_{j,j+1}^{(\ell),\gamma}(\eta)$. The same argument applies to the full-torus rates $c_{i}^{\gamma}(\sigma)$, since $|\Delta_{i}H_{\gamma}(\sigma)|$ is uniformly bounded by the $L^{1}$-normalization of the kernel. ∎
@@
The first inequality is the standard entropy inequality; see [17, Lemma 6.1.4] or [14, Section 2]. Applying it to $F=|Y|$ and using
@@
Note that the Markov generator $\mathscr{L}_{\gamma}^{N}$ is a self-adjoint non-positive operator on $L^{2}(\mu_{N,\gamma,\beta})$. We recall the standard Kipnis–Varadhan estimate for additive functionals of reversible Markov processes in the following lemma; see [18] or [17, Chapter 2, Section 1].
@@
Now we consider the fast process $(\sigma(\alpha_{\gamma}^{-1}t))_{t\geq 0}$ and its associated accelerated generator $\mathscr{A}_{N}:=\alpha_{\gamma}^{-1}\mathscr{L}_{\gamma}^{N}$ with total magnetization $M$. Note that the canonical Gibbs measure $\mu_{N,\gamma,\beta}$ is also reversible for $\mathscr{A}_{N}$. Let $\nu_{N,t}$ be the law of the process $\sigma^{N}(\alpha^{-1}_{\gamma}t)$, so that
$\nu_{N,t}(d\sigma)=f_{N,t}(\sigma)\,\mu_{N,\gamma,\beta}(d\sigma),\quad f_{N,t}\geq 0,\quad\int_{\sum_{N,M}}f_{N,t}\,d\mu_{N,\gamma,\beta}=1.$
We define the relative entropy with respect to the canonical Gibbs equilibrium by
$H_{N}(f_{N,t}\mid\mu_{N,\gamma,\beta}):=\int_{\sum_{N,M}}f_{N,t}(\sigma)\log f_{N,t}(\sigma)\,\mu_{N,\gamma,\beta}(d\sigma).$
Now we recall the standard results for entropy dissipation and entropy production rate of the Ising–Kac–Kawasaki dynamics, see e.g. [8].
@@
Let $(f_{N,t})_{t\in[0,T]}$ be the density of the Ising–Kac–Kawasaki dynamics $\sigma^{N}(\alpha^{-1}_{\gamma}t)$ with respect to $\mu_{N,\gamma,\beta}$. Then the relative entropy is absolutely continuous in time and satisfies
$\frac{d}{dt}H_{N}(f_{N,t}\mid\mu_{N,\gamma,\beta})=-4\alpha^{-1}_{\gamma}\sum_{x\in\Lambda_{N}}I_{x,x+1}^{N,\gamma}(f_{N,t})=-4\alpha^{-1}_{\gamma}I_{N}^{\gamma}(f_{N,t})$
for almost every $t\in[0,T]$. Therefore
$4\alpha^{-1}_{\gamma}\int_{0}^{T}\sum_{x\in\Lambda_{N}}I_{x,x+1}^{N,\gamma}(f_{N,t})\,dt\leq H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta}).$
Since the canonical Gibbs measure $\mu_{N,\gamma,\beta}$ is also reversible for $\mathscr{A}_{N}$, we have the following relation:
$\|W\|_{-1,\mathscr{A}_{N}}^{2}:=\sup_{f:\Sigma_{N}\to\mathbb{R}}\Bigl\{2\langle W,f\rangle_{\mu_{N,\gamma,\beta}}-\langle f,(-\mathscr{A}_{N})f\rangle_{\mu_{N,\gamma,\beta}}\Bigr\}=\alpha_{\gamma}\|W\|_{\mathcal{H}^{-1}_{N}}^{2}$
Since $0<\alpha_{\gamma}\ll 1$, and the ergodicity of $\mu_{N,\gamma,\beta}$, we show the Entropy–Kipnis–Varadhan estimate for additive functionals with the accelerated process $(\sigma^{N}(\alpha_{\gamma}^{-1}t))_{t\geq 0}$ with non-equilibrium initial laws.
Lemma 9.
Let $(\sigma^{N}(\alpha_{\gamma}^{-1}t))_{t\geq 0}$ be the Markov process generated by $\mathscr{A}_{N}$, started from the initial law $f_{N,0}\mu_{N,\gamma,\beta}$, and assume $H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})\leq C_{H}N.$ For a bounded observable $W:\Sigma_{N}\to\mathbb{R}$ with $\int W\,d\mu_{N,\gamma,\beta}=0,$ we have estimates
$\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}W(\sigma^{N}(\alpha_{\gamma}^{-1}t))\,dt\right|\leq C\,\sqrt{\alpha_{\gamma}NT}\,\|W\|_{\mathcal{H}^{-1}_{N}}.$ (12)
@@
Lemma 6 with $Y:=\int_{0}^{T}W(\sigma(\alpha_{\gamma}^{-1}t))\,dt,\mu=f_{N,0}\mu_{N,\gamma,\beta}$, and $\nu=\mu_{N,\gamma,\beta}$ yields that for every $a>0$,
$\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}W(\sigma(\alpha_{\gamma}^{-1}t))\,dt\right|\leq\frac{H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})+\log 2}{a}+\frac{1}{a}\max_{\pm}\log\mathbb{E}_{\mu_{N,\gamma,\beta}}\left[\exp\!\left(\pm a\int_{0}^{T}W(\sigma(\alpha_{\gamma}^{-1}t))\,dt\right)\right].$
This gives the entropy bound
$\mathbb{E}_{f_{N,0}}|Y|\leq\frac{H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})+\log 2}{a}+\frac{1}{a}\max_{\pm}\log\mathbb{E}_{\mu_{N,\gamma,\beta}}\bigl[e^{\pm aY}\bigr].$ (13)
@@
$\log\mathbb{E}_{\mu_{N,\gamma,\beta}}\left[\exp\!\left(\pm a\int_{0}^{T}W(\sigma(\alpha_{\gamma}^{-1}t))\,dt\right)\right]\leq C_{\mathrm{KV}}\,a^{2}T\alpha_{\gamma}\|W\|_{\mathcal{H}^{-1}_{N}}^{2}.$
Substituting this estimate into the entropy bound (13) yields
$\mathbb{E}_{f_{N,0}}|Y|\leq$ $\frac{H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})+\log 2}{a}+C_{\mathrm{KV}}aT\alpha_{\gamma}\|W\|_{\mathcal{H}^{-1}_{N}}^{2}$
$\leq$ $C\sqrt{\bigl(H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})+1\bigr)\,T\alpha_{\gamma}}\|W\|_{\mathcal{H}^{-1}_{N}}.$
@@ 3 Expansion for the discrete drift
In this section we give an expansion for the discrete drift for the rescaled field $X_{\gamma}$.
With the Ising–Kac Hamiltonian
$H_{\gamma}(\sigma)=-\frac{1}{2}\sum_{i,j\in\Lambda_{N}}\kappa_{\gamma}(i-j)\sigma_{i}\sigma_{j}=-\frac{1}{2}\sum_{j\in\Lambda_{N}}h(\sigma,j)\sigma_{j},$
@@
$X_{\gamma}(t,x)=X_{\gamma}(0,x)-\int_{0}^{t}\nabla^{*}_{\varepsilon_{\gamma}}\mathcal{J}_{\gamma}(s,x)\,ds+M_{\gamma}(t,x),$
@@
$\Bigl(\alpha_{\gamma}^{-1}\mathscr{L}_{\gamma}^{N}\langle X_{\gamma},\phi\rangle_{\gamma}\Bigr)(\sigma)=\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}\mathcal{J}_{\gamma}(\sigma;\varepsilon_{\gamma}i)\nabla_{\varepsilon_{\gamma}}\phi(\varepsilon_{\gamma}i)=\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}j_{i}(\sigma)\,\nabla_{\varepsilon_{\gamma}}\bigl(\mathcal{K}_{\gamma}^{\varepsilon}\phi\bigr)(\varepsilon_{\gamma}i)$
@@
$(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i):=\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(i-j)\phi(\varepsilon_{\gamma}j).$
Since $X_{\gamma}(t,\varepsilon_{\gamma}i)=\delta_{\gamma}^{-1}h_{\gamma}\!\left(\sigma^{N}(t/\alpha_{\gamma});i\right),\varepsilon_{\gamma}i\in\mathbb{T}$, we have
$h_{\gamma}\!\left(\sigma^{N}\left(\frac{t}{\alpha_{\gamma}}\right);i\right)-h_{\gamma}\!\left(\sigma^{N}\left(\frac{t}{\alpha_{\gamma}}\right);i+1\right)=-\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i),$
@@
with $\widetilde{R}_{i}=\frac{1}{6}d_{i}F^{(3)}(a_{i}+\theta_{i}b_{i})b_{i}^{3},\ \theta_{i}\in(0,1).$ Therefore
@@
(i) Linear part. Since $d_{i}^{2}\in\{0,4\}$, we have $\beta a_{i}\in\{0,4\kappa_{\gamma}(1)\}$, hence $F(a_{i})$ takes only two values. Using the Taylor expansion $F(u)=\frac{1}{2}-\frac{u}{4}+O(u^{2})$ at $u=0$, one obtains
@@
(ii) Nonlinear part. Since $F^{\prime}(u)=-\frac{1}{4}+O(u)$, and $d_{i}b_{i}=-d_{i}^{2}\,\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}$, we obtain the leading nonlinear contribution
@@
Note that $d_{i}^{2}(t/\alpha_{\gamma})=(\sigma_{i+1}(t/\alpha_{\gamma})-\sigma_{i}(t/\alpha_{\gamma}))^{2}$ is a accelerated process with fast time scale $\alpha_{\gamma}$. This is the term whose second-order Boltzmann–Gibbs closure produces the cubic drift in the limit.
(iii) Remainder. By construction, $j_{i}^{\mathrm{rem}}=\frac{1}{2}\big(d_{i}F^{\prime\prime}(a_{i})(b_{i})^{2}\big)(t)+\widetilde{R}_{i}^{\gamma}(t).$
@@
$\mathcal{J}_{\gamma}^{\mathrm{lin}}(t,x)$ $:=\frac{\varepsilon_{\gamma}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}\kappa_{\gamma}\!\left(\frac{x}{\varepsilon_{\gamma}}-i\right)j_{i}^{\mathrm{lin}}\left(\sigma\!\left(\frac{t}{\alpha_{\gamma}}\right)\right)$ (15)
@@
$\mathcal{J}_{\gamma}^{\mathrm{nl}}(t,x)$ $:=\frac{\varepsilon_{\gamma}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}\kappa_{\gamma}\!\left(\frac{x}{\varepsilon_{\gamma}}-i\right)\,j_{i}^{\mathrm{nl}}\left(\sigma\!\left(\frac{t}{\alpha_{\gamma}}\right)\right),$ (16)
@@
$\mathcal{J}_{\gamma}^{\mathrm{rem}}(t,x)$ $:=\frac{\varepsilon_{\gamma}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}\kappa_{\gamma}\!\left(\frac{x}{\varepsilon_{\gamma}}-i\right)j_{i}^{\mathrm{rem}}.$ (17)
We estimate the linear term and the remainder term in this section. In order to finish these estimates, We need the following deterministic estimates for the discrete gradient of the coarse field $X_{\gamma}$.
Lemma 10.
There exists a deterministic constant $C>0$, independent of $\gamma$ and $T$, such that
$\sup_{0\leq t\leq T}\sup_{x\in\Lambda_{\varepsilon_{\gamma}}}\delta_{\gamma}\varepsilon_{\gamma}\bigl|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,x)\bigr|\leq C\gamma\rightarrow 0\qquad\text{as }\gamma\downarrow 0.$ (18)
Proof.
Let $x=\varepsilon_{\gamma}k\in\Lambda_{\varepsilon_{\gamma}}$. By definition of the discrete gradient and $X_{\gamma}$,